OpenStudy (anonymous):

A company has tow electric motors consume varying amounts of power. The power consumed by each motor is a function of the time (t in minutes) for which it runs. The cost of power (in $) to run one motor is given by the function Ca(t)=t^2-2t+5. The cost of running the second motor is given by Cb(t)=3t+2. Which gives the total cost of running both motors? C(t)=3t^3-6t^2+15t C(t)=2t^2-4t+10 C(t)=t^2+t+7 C(t)=3t^3+6t^2-15t

OpenStudy (anonymous):

Cost of running both would be Cost of 1st + Cost of 2nd. So you can add both equations to find the cost of both.

OpenStudy (anonymous):

\[C_a(t)= t^2-2t+5\]\[+C_b(t) = 3t+2\]_________________ =?

OpenStudy (anonymous):

t^2+t+7?

OpenStudy (anonymous):

Correct! Great Job

OpenStudy (anonymous):

Yay! Thank you! Do you mind helping me with just three more problems?